# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v1

@@ Discussion
For Masoprocol, we were able to predict the correct target gene EGFR as the fifth prediction in settings P1 and P2. However, although EGFR is the sole target gene of Masoprocol according to the information deposited in KEGG, it is also reported to have inhibitory effects towards lipoxygenase activity [42]. Since ALOX5 is a gene that codes a lipoxygenase, treating this prediction as inaccurate should be considered incorrect. The diseases in which Masoprocol is used for its treatment include actinic keratosis [43], and it has been previously known that the mutation of TP53 is involved in the onset of this disease [44, 45]. Also, the gene that was ranked the highest in this setting, ALOX5, is reported to be one of the transcription targets of TP53 [46]. Taken together, this information suggests that TP53 is deeply related to the target gene and symptoms of Masoprocol, and therefore, although it may not be a direct target, it was highly ranked in the prediction.
@@
When we set Poziotinib as the query, EGFR and ERBB2 were correctly ranked the highest in settings P1 and P2. However, in setting G, ROS1, EML4, and ALK were also found among the highly ranked targets. Since Poziotinib was initially developed as an effective drug to treat lung cancer with HER2 mutation [47], it can be interpreted that such context was reflected in the embeddings, as ROS1 and EML4-ALK fusion genes are also characteristic mutations and therapeutic targets in lung cancer [48, 49].
@@
Selumetinib and Trametinib both target MAP2K1 and MAP2K2, and in both cases these genes were found within the top 10 predictions. Aside from these target genes, we observed BRAF and PI3KCA placed in higher rank among the predicted genes. Considering the history that these drugs were both developed for treating cancers with BRAF V600 mutation and that combination treatment with inhibition of the PI3K-AKT pathway have been explored [50, 51, 52], we assumed that such background was reflected in these results.
@@
As we explored in this section, it was suggested that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we sought to predict in this study, but also integrate broader higher-order information. In this specific analysis, predicted genes are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or those that are alternative therapeutic targets for the target disease.
@@ Conclusions
In this study, we used the embeddings learned from biological texts and performed analogy tasks for predicting drug-gene relations. We defined vectors representing these relations and showed that these vectors can accurately predict the target genes for given drugs. Furthermore, we categorized drugs and genes based on biological pathways and then defined vectors representing drug-gene relations for each pathway. Analogy computation using these vectors showed more performance improvement. In addition, we divided datasets by year and redefined the vectors representing known relations to predict unknown relations. The experimental results show that our approach can predict unknown relations to some extent.
@@ 1.2.1 Global setting by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (21), we prepare the evaluation of the analogy tasks. For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D^{y\mid I}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{D}^{y}$ and $G^{y\mid I}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{G}^{y}$. Also we define $D^{y}:=D^{y\mid(-\infty,\infty)}$ and $G^{y}:=G^{y\mid(-\infty,\infty)}$.
Similar to Eq. (11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]^{y\mid I}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y\mid I}$, and $[g]^{y\mid I}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y\mid I}$. These are formally defined as follows:
@@
In particular, we define $[d]^{y}:=[d]^{y\mid(-\infty,\infty)}$ and $[g]^{y}:=[g]^{y\mid(-\infty,\infty)}$. Given the above, we perform the analogy tasks in the following setting.
Setting Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
@@
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in D^{y}$, the set of answer genes is $[d]^{y}$, where $[d]^{y}=[d]^{y\mid(-\infty,\infty)}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
Note the following in settings Y1 and Y2:
  • In setting Y1, the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensures that predicted target genes have the relations that appeared only after year $y$.
  • In setting Y2, to evaluate the ability to predict target genes regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}^{y}$ instead of $(d,g)\in\mathcal{R}^{y\mid U_{y}}$ in Eq. (20).
BCV Ours
@@ 1.2.2 Pathway-wise setting by year
For $(d,g)\in\mathcal{R}_{p}^{y}$, we divide $\mathcal{R}_{p}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. Similar to $\mathcal{R}^{y\mid I}$ in Eq. (19), for an interval $I$, we define the set $\mathcal{R}_{p}^{y\mid I}$, which consists of all $(d,g)\in\mathcal{R}_{p}^{y}$ such that $y_{(d,g)}\in I$, as follows:
$\mathcal{R}^{y|I}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in I\}.$ (S5)
Using $L_{y}=(-\infty,y]$ and $U_{y}=(y,\infty)$, the set of drug-gene pairs that satisfies $y_{(d,g)}\leq y$ is $\mathcal{R}_{p}^{y\mid L_{y}}$, and the set of drug-gene pairs that satisfies $y<y_{(d,g)}$ is $\mathcal{R}_{p}^{y\mid U_{y}}$. By definition, $\mathcal{R}_{p}^{y\mid L_{y}}\cap\mathcal{R}_{p}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}_{p}^{y\mid L_{y}}\cup\mathcal{R}_{p}^{y\mid U_{y}}=\mathcal{R}_{p}^{y\mid(-\infty,\infty)}=\mathcal{R}_{p}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}_{p}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}_{p}^{y\mid U_{y}}$. Using the vector $\mathbf{v}_{p}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}_{p}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y}$ to $\mathbf{u}_{d}$:
@@
$\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S7)
@@
where $\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y\mid L_{y}}$. Similar to Eq. (7), Eq. (S7) defines the estimator $\hat{\mathbf{v}}_{p}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}$. For easy comparison with Eq. (5), Eq. (S7) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S8)

### v2

@@ Results
For Masoprocol, we were able to predict the correct target gene EGFR as the fifth prediction in settings P1 and P2. However, although EGFR is the sole target gene of Masoprocol according to the information deposited in KEGG, it has also been reported to inhibit lipoxygenase activity [43]. Since ALOX5 codes for a lipoxygenase, this prediction should not be considered inaccurate. The diseases in which Masoprocol is used for its treatment include actinic keratosis [44], and it has been previously known that the mutation of TP53 is involved in the onset of this disease [45, 46]. Also, the gene that was ranked the highest in this setting, ALOX5, is reported to be one of the transcription targets of TP53 [47]. Together, this information suggests that TP53 is deeply related to the target gene and symptoms of Masoprocol. Therefore, although it may not be a direct target, its high rank in the prediction is justified.
@@
When we set Poziotinib as the query, EGFR and ERBB2 were correctly ranked the highest in settings P1 and P2. However, in setting G, ROS1, EML4, and ALK were also found among the highly ranked targets. Since Poziotinib was initially developed as an effective drug to treat lung cancer with HER2 mutation [48], this context appears to be reflected in the embeddings, as ROS1 and EML4-ALK fusion genes are also characteristic mutations and therapeutic targets in lung cancer [49, 50].
@@
Selumetinib and Trametinib both target MAP2K1 and MAP2K2, and in both cases these genes were found within the top 10 predictions. Aside from these target genes, we observed BRAF and PI3KCA placed in higher rank among the predicted genes. Given the history that these drugs were both developed to treat cancers with BRAF V600 mutation and that combination treatment with inhibition of the PI3K-AKT pathway has been explored [51, 52, 53], we assumed that such background was reflected in these results.
Discussion
Mikolov et al. [54] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [55]. Given this, the performance of BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs, as presented in Table 3, is comparable to that of skip-gram in the basic analogy tasks. Also, Tshitoyan et al. [18] reported an overall accuracy of 60.1% when using analogy computation of word embedding to predict 29,046 relationships across various concepts in materials science, including element names, crystal symmetries, and magnetic properties. Our results for drug-gene relationship prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well. This demonstrates the potential of this method across different scientific disciplines. However, it’s important to note that the accuracy varies significantly across different types of relationships in the materials science study. For instance, predictions for chemical element names showed high accuracy (71.4%), while crystal structure names had lower accuracy (18.7%). This variability suggests that the effectiveness of word embedding analogies may depend on the specific type of relationship being predicted.
In Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4. It is rather surprising that simple vector addition can achieve performance comparable to large language models. Given that the GPT series is trained on vast amounts of text data, using relation vector addition is an efficient method. Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relationships is encoded in the embeddings. This also suggests that the computation of analogies may be a fundamental inference principle within large language models.
In addition, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations. The experimental results presented in Fig. 3 demonstrate that our approach can predict unknown future relations to some extent. The experiments conducted with year-split datasets are a novel attempt not found in existing research, and confirmed the effectiveness of predicting unknown future events.
@@ Discussion
In Table 4, predicted genes in the ErbB signaling pathway are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or alternative therapeutic targets for the target disease. This may suggest that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we aimed to predict in this study but also integrate broader, higher-order information.
@@ Conclusions
In this study, we used embeddings learned from biological texts and performed analogy tasks to predict drug-gene relations. We defined vectors representing these relations and showed that these vectors can accurately predict the target genes for given drugs. Additionally, we categorized drugs and genes based on biological pathways and defined vectors representing drug-gene relations for each pathway. Analogy computations with these vectors showed performance improvement. Our analogy computations demonstrated performance comparable to analogy tasks in other fields and predictions by state-of-the-art large language models, reinforcing the effectiveness of relationship prediction through simple vector addition. Moreover, the experiments with year-split datasets demonstrated that it is possible to predict unknown future relations. Not only were the predictions highly accurate, but the top genes predicted from drugs by our analogy computations were also confirmed to be reasonable from the perspective of biological expertise.
@@ Acknowledgements
Additional information
The authors declare that they have no competing interests.
@@ 1.2.1 Global setting by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20), we prepare the evaluation of the analogy tasks. Using the projection operations in Eq. (10), we define $D^{y}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y})\subset\mathcal{D}^{y}$ and $G^{y}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y})\subset\mathcal{G}^{y}$.
Similar to Eq. (11), we define $[d]^{y}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y}$, and $[g]^{y}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y}$. These are formally defined as follows:
$[d]^{y}:=\{g\mid(d,g)\in\mathcal{R}^{y}\}\subset\mathcal{G}^{y},\quad[g]^{y}:=\{d\mid(d,g)\in\mathcal{R}^{y}\}\subset\mathcal{D}^{y}.$ (S3)
Given the above, we perform the analogy tasks in the following setting.
Setting Y1.
In the analogy tasks, the set of answer genes for a query drug $d\in D^{y}$ is $[d]^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}^{y}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$. For the top-$k$ accuracy, if any of the top $k$ predictions $\hat{g}_{d}^{(1)},\ldots,\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
1.2.2 Global setting to predict unknown relations by year
Similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y\mid L_{y}}$:
$\hat{\mathbf{v}}^{y\mid L_{y}}:=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (S4)
where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$. We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq. (S4) by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20). For easy comparison with Eq. (5), Eq. (S4) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S5)
$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S6)
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq. (S4), we prepare the evaluation of the analogy tasks. For $I\in\{L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D^{y\mid I}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{D}^{y}$ and $G^{y\mid I}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{G}^{y}$.
Similar to Eq. (11), for $I\in\{L_{y},U_{y}\}$, we define $[d]^{y\mid I}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y\mid I}$, and $[g]^{y\mid I}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y\mid I}$. These are formally defined as follows:
@@ 1.2.2 Global setting to predict unknown relations by year
Given the above, we perform the analogy tasks in the following setting.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
Note the following in setting Y2:
  • In setting Y2, the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensures that predicted target genes have the relations that appeared only after year $y$.
BioConceptVec Our embeddings
@@
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 16.346 16.302
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 101.765 91.488
@@ 1.2.3 Pathway-wise setting by year
For drug-gene pairs $(d,g)\in\mathcal{R}_{p}^{y}$, we consider the analogy tasks for predicting the target genes $g$ from a drug $d$. To solve these analogy tasks, we use the relation vector $\mathbf{v}_{p}^{y}$, which represents the relation between drugs and target genes categorized in the same pathway $p$. We predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y}$ to $\mathbf{u}_{d}$:
@@
$\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S10)
@@
where $\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y}$. Similar to Eq. (7), Eq. (S10) defines the estimator $\hat{\mathbf{v}}_{p}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y}$. For easy comparison with Eq. (5), Eq. (S10) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\},$ (S11)
$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d}.$ (S12)
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y}$ in Eq. (S10), we prepare the evaluation of the analogy tasks. Using the projection operations in Eq. (10), we define $D_{p}^{y}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y})$ and $G_{p}^{y}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y})$.
Similar to Eq. (11), we define $[d]_{p}^{y}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y}$, and $[g]_{p}^{y}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y}$. These are formally defined as follows:
$[d]_{p}^{y}:=\{g\mid(d,g)\in\mathcal{R}_{p}^{y}\}\subset\mathcal{G}_{p}^{y},\quad[g]_{p}^{y}:=\{d\mid(d,g)\in\mathcal{R}_{p}^{y}\}\subset\mathcal{D}_{p}^{y}.$ (S13)
Similar to settings P1, P2, Y1, we perform the analogy tasks in the following four settings.
Setting P1Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct. In other words, for a query drug $d\in D_{p}^{y}$, the set of answer genes is $[d]_{p}^{y}$. The search space is the set of all genes $\mathcal{G}^{y}$, not limited to $\mathcal{G}_{p}^{y}$, the set of genes categorized in the pathway $p$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}^{y}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y}$, then the prediction is considered correct.
Setting P2Y1.
The gene predictions $\hat{g}_{d}$ and $\hat{g}_{d}^{(k)}$ are defined exactly the same as those in setting P1Y1, but the answer genes are defined the same as in setting Y1. That is, for the target genes that have drug-gene relations with a drug $d$, genes are considered correct regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not. In other words, for a query drug $d\in D^{y}$, the set of answer genes is $[d]^{y}$, and the prediction is considered correct if $\hat{g}_{d}\in[d]^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct. Note that the experiment is performed for $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ for each $p$.
1.2.4 Pathway-wise setting to predict unknown relations by year
For $(d,g)\in\mathcal{R}_{p}^{y}$, we define two subsets of $\mathcal{R}_{p}^{y}$ based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. Similar to $\mathcal{R}^{y\mid L_{y}}$ and $\mathcal{R}^{y\mid U_{y}}$ in Eq. (23), using $L_{y}=(-\infty,y]$ and $U_{y}=(y,\infty)$, we define the set $\mathcal{R}_{p}^{y\mid L_{y}}$ and $\mathcal{R}_{p}^{y\mid U_{y}}$ as follows:
$\mathcal{R}^{y|L_{y}}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in L_{y}\},\mathcal{R}^{y|U_{y}}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in U_{y}\}\subset\mathcal{R}_{p}^{y}.$ (S14)
The set of drug-gene pairs that satisfies $y_{(d,g)}\leq y$ is $\mathcal{R}_{p}^{y\mid L_{y}}$, and the set of drug-gene pairs that satisfies $y<y_{(d,g)}$ is $\mathcal{R}_{p}^{y\mid U_{y}}$. By definition, $\mathcal{R}_{p}^{y\mid L_{y}}\cap\mathcal{R}_{p}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y|L_{y}}_{p}\cup\mathcal{R}^{y|U_{y}}_{p}\subset\mathcal{R}_{p}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}_{p}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}_{p}^{y\mid U_{y}}$. Using the vector $\mathbf{v}_{p}^{y\mid L_{y}}$, which represents the drug-gene relations and is derived from $\mathcal{R}_{p}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$:
$\mathbf{u}_{d}+\mathbf{v}_{p}^{y\mid L_{y}}\approx\mathbf{u}_{g}.$ (S15)
Similar to equations (13) and (24), Eq. (S15) corresponds to Eq. (4). Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}_{p}^{y\mid L_{y}}$ in Eq. (S15) is defined as follows:
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}:=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S16)
where $\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y\mid L_{y}}$. Similar to Eq. (7), Eq. (S16) defines the estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}$. For easy comparison with Eq. (5), Eq. (S16) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S17)